OpenStudy (anonymous):

Solve 3x – 7y = 25 using matrices. 5x – 8y = 27 (25, 27) (27, 25) (-1, -4) (4, 1)

OpenStudy (anonymous):

\[\left[\begin{matrix}3 & -7\\5&-8\end{matrix}\right] \left(\begin{matrix}x \\ y\end{matrix}\right) = \left(\begin{matrix}25 \\ 27\end{matrix}\right)\] Now find the inverse of the first matrix and multiply.

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

what 11?

OpenStudy (anonymous):

that wat i got

OpenStudy (anonymous):

Oh, it is the determinant(?)

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

Okay, do you know how to find the inverse of a matrix?

OpenStudy (anonymous):

i make b and c negative right?

OpenStudy (anonymous):

Inverse of \[\left[\begin{matrix}a & b \\c & d\end{matrix}\right]\] is \[\frac{ 1 }{\left| A \right| } \left[\begin{matrix}d & -b \\ -c &a\end{matrix}\right]\] where /A/ is the determinant.